# Towards Combinatorial Interpretability of Neural Computation

### v1

@@ 6.4 Decoding the Circuit
Figure 22: Disentangling of networks computing the feature circuit depicted in Figure 21. In top a 2-layer network with an initial Hadamard embedding layer $C_{0}$ that is trained as before on a DNF of 8 different 2-AND clauses. The raw trained $W_{1}$ matrix is shown immediately to the left of it. Then to its left we show how one can disentangle $C_{1}$ from $W_{1}$ by multiplying $W_{1}$ by the Hadamard embedding $C_{0}$. Notice that $C_{1}$ has the usual expected structure with coding of 2P in the positive witness rows and coding for a XOR using 1N1P in the negative witness rows. Moreover, though having less 2P patterns and positive rows, it has a similar structure to the case where there is no embedding (the embedding is the identity matrix) as shown in the lower part of the figure. The bottom row shows how a symmetric random embedding is disentangled in a similar fashion.
Figure 22 shows that Cascading Disentanglement works for a minimal, small example network: the network is learning $D_{1}\approx C_{0}^{-1}$ via gradient descent as part of the learning of $W_{1}$. The feature circuit depicted in Figure 21 is captured by the matrices $W_{1}$ (second from right) and $W_{2}$ (on the left) in Figure 22, together with biases that are not shown. The middle part of the figure shows what happens when the embedding $C_{0}$ is the identity function, the top part when it is a Hadamard matrix, and the bottom when it is a symmetric random embedding. As can be seen, the trained $W_{1}$ matrix has learned the encoding of the Layer 1 feature circuit. In the case of the identity embedding in the middle part, the coding is clear. In the case of the Hadamard and random embeddings, the $W_{1}$ matrices look pretty random. However, we reveal the encoding $C_{1}$ hidden within $W_{1}$ by multiplying the first layer matrix $W_{1}$ by the Hadamard (or random) embedding $C_{0}$, because as we said, $W_{1}C_{0}\approx C_{1}D_{1}C_{0}\approx C_{1}C_{0}^{-1}C_{0}$. In other words, we undo the decoding learned by the network in $W_{1}$. This can be seen in the $W_{1}C_{0}$ matrix to the left of $W_{1}$ in the top row, where this disentanglement exposes the feature channel coding. This coding in the Hadamard and random embedding cases differs from the one for the identity embedding because SGD is affected by the embedding space, but it nevertheless is a valid feature channel coding that delivers an accurate neural network implementation of the same feature circuit.
@@
This is a good place to pause and contemplate the power of disentanglement. If one were to look at the activations of the raw $W_{1}$ matrices, before disentanglement, then the activations of neurons and the relative magintude of the weights might be quite misleading. For example, in the Hadamard embedding example, it would seem from $W_{1}$ that for Neuron 3, the weights in columns 2 and 15 are the most important (as they are of the largest magnitude), and yet we know from the disentangled matrix that these weights are just the result of the embedding space: the weights in these columns are not explaining in any way what neuron 3 is contributing to the computation. From the disentangled matrix $C_{1}$ we can tell that Neuron 3 is part of the coding for fetaures $(x_{2}\land x_{12})$ and $(x_{0}\land x_{9})$. This is the power of the combinatorial disentanglement, if we wanted to decide which weights or neurons to prune, which to apply dropout to during training and so on, doing so in the disentangled space would give us the insights that are missing when using activation space approaches.
In Step 2, to get $C_{2}$ from $W_{2}$ (and in general $C_{i}$ from $W_{i}$) we cannot simply multiply $W_{2}*C_{1}$ to get $C_{2}$ since $D_{2}$ is not the left inverse of $C_{1}$. Why? Firstly, since unlike in Step 1, there is a computation of features via bias and ReLU applied to the outputs of the multiplication by $C_{1}$. Secondly, $C_{1}$ is an $n_{1}\times m_{0}$ matrix while $D_{2}$ is an $m_{1}\times n_{1}$ matrix. An inverse of $D_{2}$ would need to be an $n_{1}\times m_{1}$ matrix. In other words, one needs to decipher the unknown $m_{1}$ features that $C_{1}$ is computing-on and compressing-from, and use them to create a good approximation of $D_{2}^{-1}$. Then in general, we conjecture that with $C_{2}$ at hand we could approximate the inverse of $D_{3}$ and get $C_{3}$ and so on in a cascading sequence of disentanglement steps, decoding the full MLP of depth $d$. Implementing this approach is a topic for future research. [^4] Note that we are proposing an algorithmic framework without having the solution algorithm, apart from some initial heuristic attempts. Skeptics might criticize us for this. We note that this has been done before, for example, in the context of public key cryptography [9] where Diffie and Hellman proposed a public key cryptography algorithm without the RSA algorithm [33] to implement it.
@@
In the above experiments we have used three different types of embeddings: identity (or no embedding), using a Hadamard matrix, and using a symmetric random embedding. In most of the statistics in Section 4 we trained on a network where the input had a Hadamard embedding layer $C_{0}$ which we then used to get $W_{1}*C_{0}$, the coding matrix in which we counted the number of patterns (as a sanity check, we also ran the same training without the Hadamard embedding, getting similar results). In all other sections we used the identity embedding.
@@
Finally, one can understand today’s prevailing approach to interpreability via the training of an autoencoder [31] using our coding-decoding view of computation. The autoencoder approach is a way of learning an approximation of $D_{i}$ for a given layer $i$. The autoencoder has a very large inner layer to which the output activations of $C_{i-1}$ are mapped to separate them and expose parts of $D_{i}$ from which the features of the layer $i-1$ feature circuit can be read. The actual autoencoder is an outside trained network that does $W_{i}*W_{i}^{T}$ as opposed to $W_{i}$ times $C_{i-1}$, the inverse of its decoding, but still there is the expansion to a larger space where monosemantic features can be learned. The difference in the combinatorial approach we propose here is to attempt to decode $C_{i-1}$ and reconstruct $D_{i}$ by analyzing the matrices rather then learning features from the activations using a separate trained network. This should not however discourage researchers from attempting a combined approach where $D_{i}$ is learned from $C_{i-1}$ via a combination of an autoencoding (or other activation based learning approachs [10]) together with a combinatorial disentanglement and decoding.

### v2

@@ 6.5 Cascading Feature Disentanglement with an Embedding Layer
Figure 22: Disentangling of networks computing the feature circuit depicted in Figure 21. In top a 2-layer network with an initial Hadamard embedding layer $C_{0}$ that is trained as before on a DNF of 8 different 2-AND clauses. The raw trained $W_{1}$ matrix is shown immediately to the left of it. Then to its left we show the disentangled $C_{1}=W_{1}C_{0}$. Notice that $C_{1}$ has the usual expected structure with coding of 2P in the positive witness rows and coding for a XOR using 1N1P in the negative witness rows. Moreover, though having less 2P patterns and positive rows, it has a similar structure as the case where there is no embedding (the embedding is the identity matrix) as shown in the middle row of the figure. The bottom row shows how a symmetric random embedding is disentangled in a similar fashion.
@@
This is a good place to pause and contemplate the power of disentanglement. The relative magnitudes of the weights in the raw $W_{1}$ matrices, before disentanglement, might be quite misleading. For example, in the Hadamard embedding example, it would seem from $W_{1}$ that for Neuron 3, the weights in columns 2 and 15 are the most important (as they are of the largest magnitude), and yet we know from the disentangled matrix that these weights are just the result of the embedding space: the weights in these columns do not explain what neuron 3 is contributing to the computation. From the disentangled matrix $C_{1}$ we can tell that Neuron 3 is part of the coding for features $(x_{2}\land x_{12})$ and $(x_{0}\land x_{9})$. This is the power of combinatorial disentanglement: if we want to decide which weights or neurons to prune, which to apply dropout to during training and so on, doing so in the disentangled space would give us the insights that are missing when using activation space approaches.
@@
In the above experiments, we have used three different types of embeddings: identity (or no embedding), using a Hadamard matrix, and using a symmetric random embedding. In most of the statistics in Section 4 we trained on a network where the input had a Hadamard embedding layer $C_{0}$ which we then used to get $W_{1}*C_{0}$, the coding matrix in which we counted the number of patterns (as a sanity check, we also ran the same training without the Hadamard embedding, getting similar results). In all other sections we used the identity embedding.
@@
Finally, one can understand today’s prevailing approach to interpreability via the training of an autoencoder [29] using our coding-decoding view of computation. The autoencoder approach is a way of learning an approximation of $D_{i}$ for a given layer $i$. The autoencoder has a very large inner layer to which the output activations of $C_{i-1}$ are mapped, along with a sparsity constraint to separate them so that the features of layer $i-1$ can be read. The difference in the combinatorial approach we propose here is to attempt to decode $C_{i-1}$ and reconstruct $D_{i}$ by analyzing the network weight matrices themselves rather then learning features from the activations using a separate trained network. As mentioned above, we think there is potential in a combined approach, for example where $D_{i}^{*}$ is learned via an autoencoder (or other activation based learning approaches [10]), and then this is used to find and interpret $C_{i}$ combinatorially. This would allow our approach to be applied to a single layer, without the need for cascading.